# A Review of Bayesian Uncertainty Quantification in Deep Probabilistic Image Segmentation

### v4

@@ Abstract
Advances in architectural design, data availability, and compute have driven remarkable progress in semantic segmentation. Yet, these models often rely on relaxed Bayesian assumptions, omitting critical uncertainty information needed for robust decision-making. The resulting reliance on point estimates has fueled interest in probabilistic segmentation, but the literature remains fragmented, marked by inconsistent terminology, evaluation practices, and theoretical ambiguities. In response, this review consolidates and contextualizes foundational concepts in uncertainty modeling, including the non-trivial task of distinguishing between epistemic and aleatoric uncertainty and examining their roles across four key downstream segmentation tasks, highlighting Active Learning as particularly promising. By unifying theory, terminology, and applications, we provide a coherent foundation for researchers and identify critical challenges, such as strong assumptions in spatial aggregation, lack of standardized benchmarks, and pitfalls in current uncertainty quantification methods. We identify trends such as the adoption of contemporary generative models, driven by advances in the broader field of generative modeling, with segmentation-specific innovation primarily in the conditioning mechanisms. Moreover, we observe growing interest in distribution- and sampling-free approaches to uncertainty estimation. We further propose directions for advancing uncertainty-aware segmentation in deep learning, including pragmatic strategies for disentangling different sources of uncertainty, novel uncertainty modeling approaches and improved Transformer-based backbones. In this way, we aim to support the development of more reliable, efficient, and interpretable segmentation models that effectively incorporate uncertainty into real-world applications.
@@ 1 Introduction
Fortunately, the merits of uncertainty quantification have been well-recognized in the field of CNN-based segmentation, especially as interpretability and reliability have become central in data-driven applications. Extensive efforts have been made to align neural network optimization with Bayesian machine learning (Blundell et al., 2015; Guo et al., 2017; Kingma & Welling, 2013; Kendall & Gal, 2017), such as learning parameter distributions rather than point estimates to capture epistemic uncertainty. Additionally, explicitly modeling the likelihood of outputs enables estimation of aleatoric uncertainty. However, in practice, these two forms of uncertainty are often entangled or even conflicting (Mucsányi et al., 2024; de Jong et al., 2024; Valdenegro-Toro & Mori, 2022; Hüllermeier & Waegeman, 2021), making the distinction difficult to define and apply. This conceptual ambiguity presents significant barriers to understanding and implementation (Kahl et al., 2024; Der Kiureghian & Ditlevsen, 2009; Kirchhof et al., 2025). At the same time, uncertainty is often treated as an auxiliary tool to enhance downstream performance, rather than as a primary modeling objective. As a result, many works lack rigorous theoretical grounding, focusing instead on empirical gains without clear definitions or principled formulations of uncertainty. Furthermore, many approaches are tailored to specific datasets or modalities, limiting generalizability. Moreover, theoretical contributions often originate from adjacent fields like Bayesian deep learning or information theory, without being properly contextualized in segmentation. This disconnect has resulted in a fragmented body of work, with inconsistent terminology, evaluation metrics, and assumptions. The resulting abundance of literature can be overwhelming, even for experienced researchers.
@@
This paper aims to explore and clarify these concepts in the context of segmentation architectures and their impact on downstream tasks. Given these challenges and the renewed interest in uncertainty (Papamarkou et al., 2024; Kahl et al., 2024; Kirchhof et al., 2025), we identify a notable absence of a comprehensive, theory-informed overview in the field. Due to its application-driven nature, most existing surveys adopt a medical perspective (Jungo & Reyes, 2019; Kwon et al., 2020), often focused on specific modalities (McCrindle et al., 2021; Jungo et al., 2020; Ng et al., 2022; Roshanzamir et al., 2023). However, the connection between theoretical foundations and their diverse applications remains underexplored. Additionally, the study by Kahl et al. (2024) introduces a valuable framework for benchmarking uncertainty disentanglement in semantic segmentation, but focuses primarily on empirical evaluation within specific tasks. In contrast, our work provides a comprehensive overview that spans the full range of established uncertainty quantification methods, unifies and contextualizes the underlying theory, and standardizes terminology and notation. This broader perspective helps clarify conceptual foundations and highlights cross-cutting insights across segmentation architectures and application domains. By the end of this paper, readers will have a clear understanding of the various forms of uncertainty, their relevance to segmentation tasks, and a comprehensive grasp of the key challenges and open research directions in the field.
@@
This review paper is structured as follows. Past work with significant impact in general image segmentation is presented in Section 2. Then, the theoretical framework and notation that govern the remainder of the paper is introduced in Section 3. The theory enables us to classify the approaches into two distinct but often related methods. That is, aleatoric and epistemic uncertainty modeling methods. The role of these concepts in image segmentation will be treated in Sections 4 and 5, respectively, which include all architectures and approaches with significant impact on the field. We then correlate these modeling methods to specific applications in Section 6. Here, we also observe for which domains uncertainty quantification has been exploited. Then, this overview will be discussed further in Section 7. This section is particularly valuable for researchers seeking to identify the most suitable model for their specific segmentation task. To encourage further research in this field, we provide future recommendations by highlighting key challenges, gaps and pitfalls in literature, concluding our review in Section 8. Figure 1 illustrates a brief overview of the sections.
@@ 2 Background
In particular, following the successful application of CNNs (LeCun et al., 1998), image segmentation experienced rapid progress driven by increasingly powerful and specialized deep architectures. Notably, the Fully Convolutional Network (FCN) (Shelhamer et al., 2014) adapted the AlexNet (Krizhevsky et al., 2012), VGG16 (Simonyan & Zisserman, 2014) and GoogLeNet (Szegedy et al., 2014) architectures to enable end-to-end semantic segmentation. Furthermore, other CNN architectures such as DeepLabv3 (Chen et al., 2017), and the MobileNetv3 (Howard et al., 2019) have also been commonly used. As the research progressed, increasing success has been observed with encoder-decoder models (Noh et al., 2015; Badrinarayanan et al., 2015; Yuan et al., 2019; Ronneberger et al., 2015). Initially developed for the medical applications, Ronneberger Ronneberger et al. (2015) introduced the U-Net, which successfully relies on residual connections between the encoding-decoding path, to preserve high-frequency details in the encoded feature maps. To this day, the U-Net is still often utilized as the default backbone model for segmentation architectures, particularly in the medical domain. In fact, reports of recent research indicate that the relatively simple U-Net and nnU-Net (Isensee et al., 2020) still outperform more contemporary and complex models (Eisenmann et al., 2023; Isensee et al., 2024).
@@ 3 Probabilistic Image Segmentation
Assuming random-variable pairs $(\mathbf{Y},\mathbf{X})\sim P_{\*Y,\*X}$ that take values in $\mathcal{Y}\in\mathbb{R}^{K\times H\times W}$ and $\mathcal{X}\in\mathbb{R}^{C\times H\times W}$, respectively, then $\mathbf{Y}$ can be considered as the ground-truth of a $K$-class segmentation task and $\mathbf{X}$ as the query image. The variables $H$, $W$ and $C$ correspond to the image height, width and channel depth, respectively. Conforming to the principle of maximum entropy, the optimal parameters given the data (i.e. posterior) subject to the chosen intermediate distributions can be inferred through Bayes Theorem as
@@
where $p(\bm{\theta})$ represents the prior belief on the parameter distribution and $p(\mathbf{y}|\mathbf{x})$ the conditional data likelihood (also commonly referred to as the evidence). After obtaining a posterior with dataset $\mathcal{D}=\{\mathbf{x}_{i},\mathbf{y}_{i}\}_{i=1}^{N}$ containing $N$ images, the predictive distribution from a new datapoint $\mathbf{x}^{*}$ can be denoted as
@@
$H[\,\mathbf{Y}|\mathbf{x}^{*},\mathcal{D}\,]=\overbrace{I[\mathbf{y},\bm{\theta}|\mathbf{x}^{*},\mathcal{D}]}^{\text{epistemic}}+\underbrace{\mathbb{E}_{q(\bm{\theta}|\mathcal{D})}[\,H[\mathbf{y}|\mathbf{x}^{*},\bm{\theta}]\,]}_{\text{aleatoric}}.$ (3)
In this way, it is theoretically possible to identify whether high-entropy predictions are due to model ignorance or statistical ambiguity inherent in the data generating process. Nonetheless, determining the nature of uncertainty is not often straightforward. For example, Hüllermeier & Waegeman (2021) stated that “by allowing the learner to change the setting, the distinction between these two types of uncertainty will be somewhat blurred". This sentiment is also shared by Der Kiureghian & Ditlevsen (2009), noting that “in one model an addressed uncertainty may be aleatory, in another model it may be epistemic". Sharing similar views, we highlight the necessity of careful analyses and possible subjective interpretation regarding the topic as we treat the realm of quantifying spatially correlated uncertainty. Especially with increasingly complex methodologies, treating the uncertainties as separate concepts is mostly theoretical (often even more philosophical) and highly non-trivial in practice.
@@ 3.1 Conventional segmentation
which simplifies the training procedure by taking a point estimate of the posterior. This approximation improves as the training data increases and the model parameter variances approach zero. As such, MLE does not include any prior knowledge on the structure of the parameter distribution but can be achieved through Bayesian Maximum A Posteriori (MAP) estimation with
@@
For example, assuming Gaussian or Laplacian priors can be a consequence of regularizing the $L_{2}$ norm (also known as ridge regression or weight decay) or $L_{1}$ norm of $\bm{\theta}$, respectively (Figueiredo, 2001; Kaban, 2007). Additionally, the output of such models can be interpreted as a probability distribution. For instance, consider spatial image dimensions $D$ and a CNN model $f_{\bm{\theta}}:\mathbb{R}^{C\times D}\rightarrow\mathbb{R}^{K\times D}$ with parameters $\bm{\theta}$, such that $\mathbf{a}=f_{\theta}(\mathbf{x})$. Then, the output of such model can be regarded as parameters of a Probability Mass Function (PMF) through
@@ 4 Aleatoric uncertainty
with Dirac-delta function $\delta$, and mapping $F:\mathcal{X}\rightarrow\mathcal{Y}$. This relationship is characterized by the ambiguity in $\mathbf{X}$ and is inherently probabilistic, due to various reasons such as noise in the data (occlusions, sensor noise, insufficient resolution, etc.) or variability within a class (e.g. not all cats have tails). Hence, the observance of substantial aleatoric uncertainty can in some cases be inevitable, but may also signal the need for higher-quality data acquisition or shifting to another modality. The possible input dependency of the uncertainty develops into further categorization of either heteroscedastic (dependent) or homoscedastic (independent) aleatoric uncertainty. In most practical scenarios, aleatoric uncertainty quantification methods encompass both types and assume a parameterized likelihood function $p(\mathbf{Y}|\mathbf{X},\bm{\theta})$ as a direct reflection of $p(\mathbf{Y}|\mathbf{X})$. For example, it is possible to model a distribution parameterized by the output of a CNN (Section 4.1). Also, conditional generative models are used to learn the data distribution through latent variables. For the purpose of taxonomy, we categorize models based on the location of these latent (unobserved) variables: those introduced near the output are discussed in Section 4.1, while lower-dimensional latents embedded deeper within the architecture are covered in Section 4.2. While their theoretical formulations are largely similar, their practical implications can differ significantly.
@@ 4.1.1 Independence
As discussed earlier, neural network predictions are often normalized with SoftMax activation in order to interpret the confidence values as parameters of a probability mass distribution. However, these values can only be interpreted as probabilities after proper validation, which is referred to as model calibration. Here, it is measured whether the empirical accuracy of a model approximately equals the provided class confidence $c_{k}$ for class $k$, i.e. $P(Y=k\,|\,c_{k})=c_{k}$. Calibration is typically visualized with a reliability diagram, where miscalibration and under-/over-confidence can be assessed by inspecting the deviation from the graph diagonal (Figure 4). Furthermore, different methods can be used to quantify calibration, although each may introduce its own biases. A fairly straightforward metric, the Expected Calibration Error (ECE), determines the normalized distance between accuracy (acc) and confidence (conf) bins as
@@ 4.1.2 Spatial correlation
Assuming pixel independence inhibits spatially coherent uncertainty quantification, which is a key factor for segmentation. Introducing spatial correlation can be achieved through an autoregressive approach. For instance, we can rephrase Equation (7) to
@@
Monteiro et al. (2020) propose the Stochastic Segmentation Network (SSN), which models the output logits as a multivariate normal distribution, parameterized by the neural networks $f_{\bm{\theta}}^{\mu}$ and $f_{\bm{\theta}}^{\Sigma}$ as
@@
where the covariance matrix has a low-rank structure $\bm{\Sigma}=\mathbf{P}\mathbf{P}^{T}+\bm{\Lambda}$, with $\mathbf{P}$ having dimensionality $((K\times D)\times R)$, $R$ being a hyperparameter that controls the parameterization rank and $\bm{\Lambda}$ representing a diagonal matrix. The low-rank assumption results in a more structured and expressive distribution, while retaining reasonable efficiency. Since, the SoftMax transform pertains an intractable integral, Monte Carlo sampling is employed. SSNs can theoretically be augmented as an additional layer to any pretrained CNN (see Figure 6).

### v5

@@ Abstract
Advances in architectural design, data availability, and compute have driven remarkable progress in semantic segmentation. Yet, these models often rely on relaxed Bayesian assumptions, omitting critical uncertainty information needed for robust decision-making. The resulting reliance on point estimates has fueled interest in probabilistic segmentation, but the literature remains fragmented. In response, this review consolidates and contextualizes foundational concepts in uncertainty modeling, including the non-trivial task of distinguishing between epistemic and aleatoric uncertainty and examining their roles across four key downstream segmentation tasks, highlighting Active Learning as particularly promising. By unifying theory, terminology, and applications, we provide a coherent foundation for researchers and identify critical challenges, such as strong assumptions in spatial aggregation, lack of standardized benchmarks, and pitfalls in current uncertainty quantification methods. We identify trends such as the adoption of contemporary generative models, driven by advances in the broader field of generative modeling, with segmentation-specific innovation primarily in the conditioning mechanisms. Moreover, we observe growing interest in distribution- and sampling-free approaches to uncertainty estimation. We further propose directions for advancing uncertainty-aware segmentation in deep learning, including pragmatic strategies for disentangling different sources of uncertainty, novel uncertainty modeling approaches and improved Transformer-based backbones. In this way, we aim to support the development of more reliable, efficient, and interpretable segmentation models that effectively incorporate uncertainty into real-world applications.
@@ 1 Introduction
Fortunately, the merits of uncertainty quantification have been well-recognized in the field of CNN-based segmentation, especially as interpretability and reliability have become central in data-driven applications. Extensive efforts have been made to align neural network optimization with Bayesian machine learning (Blundell et al., 2015; Guo et al., 2017; Kingma & Welling, 2013; Kendall & Gal, 2017), such as learning parameter distributions rather than point estimates to capture epistemic uncertainty. Additionally, explicitly modeling the likelihood of outputs enables estimation of aleatoric uncertainty. However, in practice, these two forms of uncertainty are often entangled or even conflicting (Mucsányi et al., 2024; de Jong et al., 2024; Valdenegro-Toro & Mori, 2022; Hüllermeier & Waegeman, 2021), making the distinction difficult to define and apply. This conceptual ambiguity presents significant barriers to understanding and implementation (Kahl et al., 2024; Der Kiureghian & Ditlevsen, 2009; Kirchhof et al., 2025). At the same time, uncertainty is often treated as an auxiliary tool to enhance downstream performance, rather than as a primary modeling objective. As a result, many works lack rigorous theoretical grounding, focusing instead on empirical gains without clear definitions or principled formulations of uncertainty. Furthermore, many approaches are tailored to specific datasets or modalities, limiting generalizability. Moreover, theoretical contributions often originate from adjacent fields like Bayesian deep learning or information theory, and can be insufficiently contextualized within the domain of segmentation. This disconnect has resulted in a fragmented body of work, with inconsistent terminology, evaluation metrics, and assumptions. The resulting abundance of literature can be overwhelming, even for experienced researchers.
@@
This paper aims to explore and clarify these concepts in the context of segmentation architectures and their impact on downstream tasks. Given these challenges and the renewed interest in uncertainty (Papamarkou et al., 2024; Kirchhof et al., 2025), we note that a comprehensive overview in the field remains limited. Due to its application-driven nature, most existing surveys adopt a medical perspective (Jungo & Reyes, 2019; Kwon et al., 2020), often focused on specific modalities (McCrindle et al., 2021; Jungo et al., 2020; Ng et al., 2022; Roshanzamir et al., 2023). However, the connection between theoretical foundations and their diverse applications remains underexplored. Additionally, the study by Kahl et al. (2024) introduces a valuable framework for benchmarking uncertainty disentanglement in semantic segmentation, but focuses primarily on empirical evaluation within specific tasks. In contrast, our work provides a comprehensive overview that spans the full range of established uncertainty quantification methods, unifies and contextualizes the underlying theory, and standardizes terminology and notation. This broader perspective helps clarify conceptual foundations and highlights cross-cutting insights across segmentation architectures and application domains. By the end of this paper, readers will have a clear understanding of the various forms of uncertainty, their relevance to segmentation tasks, and a comprehensive grasp of the key challenges and open research directions in the field.
@@
This review paper is structured as follows. Past work with significant impact in general image segmentation is presented in Section 2. Then, the theoretical framework and notation that govern the remainder of the paper are introduced in Section 3. The theory enables us to classify the approaches into two distinct but often related methods. That is, methods targeting either aleatoric or epistemic uncertainty, which can be distinguished by either modeling the feature or parameter distribution(s). The role of these concepts in image segmentation are treated in Sections 4 and 5, respectively, which include all architectures and approaches with significant impact on the field. We then correlate these modeling methods to specific applications in Section 6. Here, we also observe for which domains uncertainty quantification has been exploited. Then, this overview is discussed further in Section 7. This section is particularly valuable for researchers seeking to identify the most suitable model for their specific segmentation task. To encourage further research in this field, we provide future recommendations by highlighting key challenges, gaps and pitfalls in literature, concluding our review in Section 8. Figure 1 illustrates a brief overview of the sections.
@@ 2 Background
In particular, following the successful application of CNNs (LeCun et al., 1998), image segmentation experienced rapid progress driven by increasingly powerful and specialized deep architectures. Notably, the Fully Convolutional Network (FCN) (Shelhamer et al., 2014) adapted the AlexNet (Krizhevsky et al., 2012), VGG16 (Simonyan & Zisserman, 2014) and GoogLeNet (Szegedy et al., 2014) architectures to enable end-to-end semantic segmentation. Furthermore, other CNN architectures such as DeepLabv3 (Chen et al., 2017), and the MobileNetv3 (Howard et al., 2019) have also been commonly used. As the research progressed, increasing success has been observed with encoder-decoder models (Noh et al., 2015; Badrinarayanan et al., 2015; Yuan et al., 2019; Ronneberger et al., 2015). Initially developed for the medical applications, Ronneberger Ronneberger et al. (2015) introduced the U-Net, which successfully relies on residual connections between the encoding-decoding path, to preserve high-frequency details in the encoded feature maps. To this day, the U-Net is still often utilized as the default backbone model for many semantic segmentation and even general image generation architectures (Ho et al., 2020; Song et al., 2020), particularly in the medical domain. In fact, reports of recent research indicate that the relatively simple U-Net (Isensee et al., 2020) still outperform more contemporary and complex models (Eisenmann et al., 2023; Isensee et al., 2024).
@@ 3 Probabilistic Image Segmentation
Assuming random-variable pairs $(\mathbf{Y},\mathbf{X})\sim P_{\*Y,\*X}$ that take values in $\mathcal{Y}\in\mathbb{Z}^{K\times H\times W}$ and $\mathcal{X}\in\mathbb{R}^{C\times H\times W}$, respectively, then instance $\mathbf{y}$ can be considered as the ground-truth of a $K$-class segmentation task and instance $\mathbf{x}$ as the query image. The variables $H$, $W$ and $C$ correspond to the image height, width and channel depth, respectively. Conforming to the principle of maximum entropy, the optimal parameters given the data (i.e. posterior) subject to the chosen intermediate distributions can be inferred through Bayes Theorem as
@@
assuming independence of $\mathbf{x}$ from $\bm{\theta}$, and where $p(\bm{\theta})$ represents the prior belief on the parameter distribution and $p(\mathbf{y}|\mathbf{x})$ the conditional data likelihood (also commonly referred to as the evidence). After obtaining a posterior with dataset $\mathcal{D}=\{\mathbf{x}_{i},\mathbf{y}_{i}\}_{i=1}^{N}$ containing $N$ images, the predictive distribution from a new datapoint $\mathbf{x}^{*}$ can be denoted as
@@
$H[\,\mathbf{Y}|\mathbf{x}^{*},\mathcal{D}\,]=\overbrace{I[\mathbf{y},\bm{\theta}|\mathbf{x}^{*},\mathcal{D}]}^{\text{epistemic}}+\underbrace{\mathbb{E}_{q(\bm{\theta}|\mathcal{D})}[\,H[\mathbf{y}|\mathbf{x}^{*},\bm{\theta}]\,]}_{\text{aleatoric}},$ (3)
where $I$ represents the mutual information. In this way, it is theoretically possible to identify whether high-entropy predictions are due to model ignorance or statistical ambiguity inherent in the data generating process. Nonetheless, determining the nature of uncertainty is not often straightforward. For example, Hüllermeier & Waegeman (2021) stated that “by allowing the learner to change the setting, the distinction between these two types of uncertainty will be somewhat blurred". This sentiment is also shared by Der Kiureghian & Ditlevsen (2009), noting that “in one model an addressed uncertainty may be aleatory, in another model it may be epistemic". Sharing similar views, we highlight the necessity of careful analyses and possible subjective interpretation regarding the topic as we treat the realm of quantifying spatially correlated uncertainty. Especially with increasingly complex methodologies, treating the uncertainties as separate concepts is mostly theoretical (often even more philosophical) and highly non-trivial in practice.
@@ 3.1 Conventional segmentation
which simplifies the training procedure by taking a point estimate of the posterior. This approximation improves as the training data increases and the model parameter variances approach zero. As such, MLE does not include any prior knowledge on the structure of the parameter distribution but can be achieved through Maximum A Posteriori (MAP) estimation with
@@
For example, assuming Gaussian or Laplacian priors leads to regularizing the $L_{2}$ norm (also known as ridge regression or weight decay) or $L_{1}$ norm of $\bm{\theta}$, respectively (Figueiredo, 2001; Kaban, 2007). To model $p(\mathbf{y}|\mathbf{x},\bm{\theta})$, we make use of function $f_{\bm{\theta}}:\mathbb{R}^{C\times D}\rightarrow\mathbb{R}^{K\times D}$ that infers the parameters of a Probability Mass Function (PMF). For instance, consider spatial image dimensions $D$ and a CNN with $\mathbf{a}=f_{\theta}(\mathbf{x})$. Then, we can write
@@ 4 Feature modeling
with Dirac-delta function $\delta$, and mapping $F:\mathcal{X}\rightarrow\mathcal{Y}$. This relationship is characterized by the ambiguity in $\mathbf{X}$ and is inherently probabilistic, due to various reasons such as noise in the data (occlusions, sensor noise, insufficient resolution, etc.) or variability within a class (e.g. not all cats have tails). Hence, the observance of substantial aleatoric uncertainty can in some cases be inevitable, but may also signal the need for higher-quality data acquisition or shifting to another modality. The possible input dependency of the uncertainty develops into further categorization of either heteroscedastic (dependent) or homoscedastic (independent) aleatoric uncertainty. In most practical scenarios, aleatoric uncertainty modeling methods encompass both types and assume a parameterized likelihood function $p(\mathbf{Y}|\mathbf{X},\bm{\theta})$ as a direct reflection of $p(\mathbf{Y}|\mathbf{X})$. For example, it is possible to model a distribution parameterized by the output of a CNN. Also, conditional generative models are used to learn the data distribution through so-called “latent” (i.e. unobserved) variables. For the purpose of taxonomy, we categorize models based on the location of these latent variables: those introduced near the output are discussed in Section 4.1, while lower-dimensional latents embedded deeper within the architecture are covered in Section 4.2. While their theoretical formulations are largely similar, their practical implications can differ significantly.
@@ 4.1.1 Independence
As discussed earlier, neural network predictions are often normalized with SoftMax activation in order to interpret the confidence values as parameters of a probability mass distribution, but rerely reflect the true probabilities in modern neural networks Guo et al. (2017). Thus, interpreting the confidences as true probabilities is often only justified after proper validation, which is referred to as model calibration. Here, it is measured whether the empirical accuracy of a model approximately equals the provided class confidence $c_{k}$ for class $k$, i.e. $P(Y=k\,|\,c_{k})=c_{k}$. Calibration is typically visualized with a reliability diagram, where miscalibration and under-/over-confidence can be assessed by inspecting the deviation from the graph diagonal (Figure 4). Furthermore, different methods can be used to quantify calibration, although each may introduce its own biases. A fairly straightforward metric, the Expected Calibration Error (ECE), determines the normalized distance between accuracy (acc) and confidence (conf) bins as
@@ 4.1.2 Spatial correlation
Assuming pixel independence inhibits spatial coherence in uncertainty quantification, which is a key factor for segmentation. Introducing spatial correlation can be achieved through an autoregressive approach. For instance, we can rephrase Equation (7) to
@@
Monteiro et al. (2020) propose the Stochastic Segmentation Network (SSN), which models the output logits as a multivariate normal distribution, parameterized by the neural networks $f_{\bm{\theta}}^{\mu}$ and $f_{\bm{\theta}}^{\Sigma}$. Given the output features $f_{\bm{\theta}}(\mathbf{x})=\mathbf{a}$ of a deterministic model, we can denote the logits distribution as
@@
where the covariance matrix has a low-rank structure $\bm{\Sigma}=\mathbf{P}\mathbf{P}^{T}+\bm{\Lambda}$, with $\mathbf{P}$ having dimensionality $((K\times D)\times R)$, with $R$ being a hyperparameter that controls the parameterization rank and $\bm{\Lambda}$ representing a diagonal matrix. The low-rank assumption results in a more structured distribution, while retaining reasonable efficiency. Monte Carlo sampling is used to generate predictions, which are then mapped to categorical values using the SoftMax activation. SSNs can theoretically be augmented to any pretrained CNNs as an additional layer (see Figure 6).